OpenStudy (anonymous):

Solve for k 6k-2z=12

OpenStudy (jim766):

6k - 2z = 12 need to get the K by it self + 2z + 2z use opp operations to get rid of the -2z

OpenStudy (jim766):

6k - 2z = 12 need to get the K by it self + 2z + 2z use opp operations to get rid of the -2z 6k = 12 + 2z now what?

OpenStudy (anonymous):

Devide by 6?

OpenStudy (jim766):

right.... 6k = 12 + 2z ___ ________ 6 6

OpenStudy (jim766):

6k = 12 + 2z ___ ________ 6 6 k = 12 + 2z ________ 6

OpenStudy (jim766):

now you need to simplify the right side

OpenStudy (anonymous):

Wouldn't it be k=1/3z+2?

OpenStudy (jim766):

12 + 2z 12 2z ------- -- + -- 6 6 6 do them serpatelely

OpenStudy (anonymous):

2 + 1/3 z

OpenStudy (jim766):

2 is right.... 1/3 is right... but the z started out on the top...and you didnt do anything with it, so it should stay on the top 2 + z/3

OpenStudy (anonymous):

The book says the answer is K=1/3 (z+6)

OpenStudy (jim766):

yeah , it means the same thing.... 1/3(z + 6) distribute 1/3(z) + 2 z/3 + 2

OpenStudy (anonymous):

Okay thank you